OpenStudy (anonymous):

Can someone please check my answers? 1. For the function ƒ(x) = 10x2 + x, evaluate and simplify the expression: f(a+h)-f(a) / h I got f(a+h)-f(a) / h = 20a+1 2. Find ƒ−1 for the function f(x) = 23-3x/6x I got f^-1 (x) = 23/6x-3

OpenStudy (anonymous):

only f(a) is divided by h?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did you arrive at 20a+1?

OpenStudy (anonymous):

to be honest i'm not really sure, I asked someone how to do it a while ago but I forgot how to do it

OpenStudy (anonymous):

when I got the answer they told me I was right but I wanted to double check

OpenStudy (anonymous):

I'm getting 10(a+h)^2+(a+h)-(10a^2+a)/h

OpenStudy (anonymous):

can you tell me how your getting it so I can understand?

OpenStudy (anonymous):

sure

geerky42 (geerky42):

Pretty sure it's \(\dfrac{f(a+h)-f(a)}{h}\) This is sorta preparation for calculus lol.

OpenStudy (anonymous):

Basically, the way i understand f(a+h) is 'replace all x in function f(x) with a+h. So, that's the first thing I did. Which gets me 10(a+h)^2+(a+h) Then, I subtract f(a) (which is the same as f(x), instead we use a in place of x) Which makes 10(a+h)^2+(a+h) - 10a^2+a The last part needs to be divided by h, which makes: \[10(a+h)^2+(a+h)- \frac{ 10a^2+a }{ h }\] Which can't really be simplified any further.

OpenStudy (anonymous):

@geerky42 that one makes more sense to me as well, that's why I asked.

OpenStudy (anonymous):

ohh okay thank you zimmah

OpenStudy (anonymous):

how about the 2nd question?

OpenStudy (anonymous):

It doesn't look right, but I don't know what it is supposed to be

OpenStudy (anonymous):

@geerky42 could you help him with the second question here, I'm stuck there.

OpenStudy (anonymous):

are you sure you didn't make a typo there? It's really f(x) = 23-3x/6x ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Maybe this link can help you: Scroll to the lightblue box http://tutorial.math.lamar.edu/Classes/CalcI/InverseFunctions.aspx

OpenStudy (anonymous):

okay thank you